# NeuralRemaster: Phase-Preserving Diffusion for Structure-Aligned Generation

### v1

@@ Abstract
Input               FLUX-Kontext      QWen-Edit                Ours
@@ 1 Introduction
Although these tasks are conceptually easier than generating an image from scratch, existing solutions are unnecessarily complex. Methods such as ControlNet [zhang2023adding], T2I-Adapter [mou2024t2i], and related variants attach auxiliary encoders, or adapter branches to inject structural input into the model. While effective, this introduces additional parameters and computational cost, paradoxically making structure-aligned generation harder than it should be.
We argue that this inefficiency stems not from the network architecture, but from the diffusion process itself. The forward diffusion process injects Gaussian noise, which destroys both the magnitude and phase components in the frequency domain. Classical signal processing [oppenheim1981importance, ruderman1994statistics, wang2005translation], however, tells us that phase encodes structure while magnitude encodes texture. Destroying the phase means destroying the very spatial coherence that structure-aligned generation depends on, forcing the model to reconstruct structure from scratch.
@@ 2 Related Work
Diffusion Models. Diffusion models have become a dominant paradigm for generative modeling, capable of representing complex data distributions with remarkable fidelity [ho2020denoising, karras2022elucidating]. They progressively corrupt data into Gaussian noise through a forward diffusion process, then learn to invert this process via iterative denoising. This framework has demonstrated state-of-the-art performance across diverse domains, including image, video, and audio generation [saharia2022photorealistic, ramesh2022hierarchical, balaji2022ediff, chen2023videocrafter1, ho2022video, popov2021grad, kong2020diffwave], as well as reinforcement learning [janner2022planning, wang2022diffusion, psenka2023learning] and robotics [ajay2022conditional, urain2023se, chi2023diffusion].
Frequency-Domain Manipulation for Diffusion. Recent work has explored frequency domain operations for diffusion models. [rissanen2023inverse] proved that diffusion’s optimal denoiser operates as a cascade of Wiener filters with growing bandwidths, naturally recovering low-frequency structure before progressively adding high-frequency details. [crabbe2024timeseries] established that frequency-domain diffusion naturally handles dual processes through mirrored Brownian motions.[yu2025dmfft] studied the effect of guidance scales at the magnitude and phase spectrum. FreeDiff [liu2024freediff] achieved universal image editing through progressive frequency truncation with timestep-dependent schedules. MASF [gao2024masf] proposed to use a frequency-dependent moving average during sampling. [gao2024fbsdiff] proposes to guide the text-driven sampling trajectory by transplanting certain DCT frequency bands from diffusion features along the reconstruction trajectory for spatially controllable generation. [gao2025ptdiffusion] proposes to generate optical illusion images using phase interpolation based on DDIM inversion. While these approaches modify the diffusion sampling process to achieve the desired behavior, $\phi$-PD introduces minimal changes to existing diffusion frameworks and preserves the original sampling dynamics.
@@
Structure-Aligned Generation with Diffusion. Most existing methods achieve structure-aligned generation by modifying the network architecture and introducing additional adaptation components. ControlNet [zhang2023adding] copies the entire U-Net encoder into a trainable encoder branch, which adds significant computation overhead. T2I-Adapter [mou2024t2i] reduces computation overhead using a lightweight adapter module but sacrifices control precision. Uni-ControlNet [zhao2023unicontrolnet] enables simultaneous utilization of multiple local controls by training two adapters. OmniControl [tan2024ominicontrol2] integrates image conditions into Diffusion Transformer (DiT) architectures with only 0.1% additional parameters by re-using the VAE and transformer blocks of the base model. ControlNeXt [peng2024controlnext] uses a lightweight convolutional module to inject control signals, and directly finetune selective parameters of the base model to reduce training costs and latency increase. SCEdit [jiang2024scedit] proposes an efficient finetuning framework that edits skip connections using a lightweight module. NanoControl [huang2024nanocontrol] aims to achieve efficient control with a LoRA-style control module. The above methods all rely on an additional module to incorporate the control signal, though some are more lightweight than others. CosmosTransfer [Alhaija2025CosmosTransfer1CW] achieves multi-modal control by combining multiple ControlNet branches, demonstrating promising applications on physical tasks; however, multiple branches introduce significant computation overhead. In contrast, $\phi$-PD does not introduce any computation overhead or additional parameters while enabling universal spatial control.
@@
Training-Free Guidance Methods. Recently, several training-free methods have been developed. [yu2023freedom] introduced FreeDoM, which leverages off-the-shelf pre-trained networks to construct time-independent energy functions that guide generation. [couairon2023zeroshot] proposed ZestGuide for zero-shot spatial layout conditioning, utilizing implicit segmentation maps extracted from cross-attention layers to align generation with input masks. [mo2024freecontrol] presented FreeControl, a training-free approach that enforces structure guidance with the base model feature extracted from the control signal. Although these methods avoid training cost, they introduce additional overhead at test time, either an external model, DDIM inversion, or multiple inferences of the base model. In contrast, $\phi$-PD can achieve training-free spatial control without any additional inference time overhead.
@@ 3.1 Frequency Domain Fundamentals
Phase-Magnitude Separation in Signal Processing. Foundational work by Oppenheim et al. [oppenheim1981importance] shows that phase primarily determines spatial structure, while magnitude largely controls texture statistics. Mixing phase and magnitude from different images produces reconstructions whose spatial layout follows the source of the phase, not magnitude (see Figure 3). This observation motivates our approach: if diffusion destroys phase, it destroys spatial geometry; if we preserve phase, we preserve structure.
@@ 3.2 Phase-Preserving Diffusion
or sampled from Rayleigh distribution directly [GoodmanStatisticalOptics]:
@@ 3.3 Training Objective
The flow matching objective learns a vector field that transports the structured noise distribution to the target image distribution. During training, given a target image $I$ and a structured noise $\hat{\epsilon}$, and a timestep $t\in[0,1]$, an intermediate image $x_{t}$ is obtained using a linear combination between $I$ and $\hat{\epsilon}$ following Rectified Flows [liu2022flow]:
@@
In DDPMs [ho2020denoising], data $x_{0}$ is gradually corrupted by Gaussian noise:
@@ 4.1.1 Datasets
ImageNetR is a test set proposed by [pnp], including 29 images of various objects and styles. While the original dataset provides prompts, these are generic image editing prompts. Since our work primarily focuses on re-rendering, we keep the editing prompts with style hints in the original dataset and added additional style prompts, resulting in a total of 8 prompts for each image. This benchmark assesses stylized re-rendering and structure preservation.
CARLA is an open-source driving simulator [dosovitskiy2017carla]. We collect 5.5 hours of driving videos from CARLA Town 4 using the simulator’s default autopilot. We then split the videos into 25 second clips and annotate a caption for each clip. For simulation enhancement, we use these captions combined with the style hint “A photorealistic video of driving”. We evaluate the effectiveness of sim-to-real transfer by testing the CARLA-trained planner on Waymo’s WOD-E2E [xu2025wode2e] validation set.
@@ 4.2 Results
Figure 9: Video re-rendering results for simulation enhancement. From top to bottom are: input, Cosmos-Transfer2.5 [Alhaija2025CosmosTransfer1CW], ours.
@@ Appendix A Additional Implementation Details
We implemented $\phi$-PD based on three models: SD 1.5 [Rombach_2022_CVPR], FLUX-dev [flux2024] and Wan2.2-14b [wan2025]. We use the implementation of these models from DiffSynthStudio [^2] https://github.com/modelscope/DiffSynth-Studio This section describes the additional implementation details for the experiment with each model.
@@ A.1 Training and Inference Details
We start from the officially released checkpoints of SD 1.5 [Rombach_2022_CVPR], FLUX-dev [flux2024], and Wan2.2-14B [wan2025], and finetune each model with phase-preserving noise. For SD 1.5, we experiment with both full finetuning and LoRA finetuning, while for FLUX-dev and Wan2.2-14B we use LoRA finetuning due to computational constraints. At inference time, for Wan2.2-14B we adopt the 4-step LoRA from LightX2V [^3] https://huggingface.co/lightx2v/Wan2.2-Lightning and apply it directly on top of our finetuned LoRA weights to accelerate sampling.

### v2

@@ Abstract
Input  e              FLUX-Kontext      QWen-Edit                Ours
@@ 1 Introduction
Although these tasks are conceptually easier than generating an image from scratch, existing solutions are unnecessarily complex. Methods such as ControlNet [42], T2I-Adapter [21], and related variants attach auxiliary encoders, or adapter branches to inject structural input into the model. While effective, this introduces additional parameters and computational cost, paradoxically making structure-aligned generation harder than it should be.
We argue that this inefficiency stems not from the network architecture, but from the diffusion process itself. The forward diffusion process injects Gaussian noise, which destroys both the magnitude and phase components in the frequency domain. Classical signal processing [23, 30, 37], however, tells us that phase encodes structure while magnitude encodes texture. Destroying the phase means destroying the very spatial coherence that structure-aligned generation depends on, forcing the model to reconstruct structure from scratch.
@@ 2 Related Work
Diffusion Models. Diffusion models have become a dominant paradigm for generative modeling, capable of representing complex data distributions with remarkable fidelity [11, 16]. They progressively corrupt data into Gaussian noise through a forward diffusion process, then learn to invert this process via iterative denoising. This framework has demonstrated state-of-the-art performance across diverse domains, including image, video, and audio generation [31, 28, 3, 4, 12, 25, 17], as well as reinforcement learning [14, 36, 26] and robotics [1, 35, 5].
Frequency-Domain Manipulation for Diffusion. Recent work has explored frequency domain operations for diffusion models. [7] argues that diffusion models of images perform approximate autoregression in the frequency domain. [40] shows that modifying the UNet frequency domain features significantly improves the generating quality for image or video generation. FreeDiff [38] introduces a fine-tuning free approach for image editing that employs progressive frequency truncation to refine the guidance of diffusion models. [27] proposed to use a frequency-dependent moving average during sampling. [2] proposes a training-free approach for image inpainting that optimizes the initial seed noise in the spectral domain. [9] proposes to generate optical illusion images using phase interpolation based on DDIM inversion. While these approaches modify the diffusion sampling process to achieve the desired behavior, $\phi$-PD introduces minimal changes to existing diffusion frameworks and preserves the original sampling dynamics.
@@
Structure-Aligned Generation with Diffusion. Most existing methods achieve structure-aligned generation by modifying the network architecture and introducing additional adaptation components. ControlNet [42] copies the entire U-Net encoder into a trainable encoder branch, which adds significant computation overhead. T2I-Adapter [21] reduces computation overhead using a lightweight adapter module but sacrifices control precision. Uni-ControlNet [43] enables simultaneous utilization of multiple local controls by training two adapters. OmniControl [32] integrates image conditions into Diffusion Transformer (DiT) architectures with only 0.1% additional parameters by re-using the VAE and transformer blocks of the base model. ControlNeXt [24] uses a lightweight convolutional module to inject control signals, and directly finetune selective parameters of the base model to reduce training costs and latency increase. SCEdit [15] proposes an efficient finetuning framework that edits skip connections using a lightweight module. NanoControl [13] aims to achieve efficient control with a LoRA-style control module. The above methods all rely on an additional module to incorporate the control signal, though some are more lightweight than others. CosmosTransfer [22] achieves multi-modal control by combining multiple ControlNet branches, demonstrating promising applications on physical tasks; however, multiple branches introduce significant computation overhead. In contrast, $\phi$-PD does not introduce any computation overhead or additional parameters while enabling universal spatial control.
@@
Training-Free Guidance Methods. Recently, several training-free methods have been developed. [41] introduced FreeDoM, which leverages off-the-shelf pre-trained networks to construct time-independent energy functions that guide generation. [6] proposed ZestGuide for zero-shot spatial layout conditioning, utilizing implicit segmentation maps extracted from cross-attention layers to align generation with input masks. [20] presented FreeControl, a training-free approach that enforces structure guidance with the base model feature extracted from the control signal. Although these methods avoid training cost, they introduce additional overhead at test time, either an external model, DDIM inversion, or multiple inferences of the base model. In contrast, $\phi$-PD can achieve training-free spatial control without any additional inference time overhead.
@@ 3.1 Frequency Domain Fundamentals
Phase-Magnitude Separation in Signal Processing. Foundational work by Oppenheim et al. [23] shows that phase primarily determines spatial structure, while magnitude largely controls texture statistics. Mixing phase and magnitude from different images produces reconstructions whose spatial layout follows the source of the phase, not magnitude (see Figure 3). This observation motivates our approach: if diffusion destroys phase, it destroys spatial geometry; if we preserve phase, we preserve structure.
@@ 3.2 Phase-Preserving Diffusion
or sampled from Rayleigh distribution directly [10]:
@@ 3.3 Training Objective
The flow matching objective learns a vector field that transports the structured noise distribution to the target image distribution. During training, given a target image $I$ and a structured noise $\hat{\epsilon}$, and a timestep $t\in[0,1]$, an intermediate image $x_{t}$ is obtained using a linear combination between $I$ and $\hat{\epsilon}$ following Rectified Flows [19]:
@@
In DDPMs [11], data $x_{0}$ is gradually corrupted by Gaussian noise:
@@ 4.1.1 Datasets
ImageNetR is a test set proposed by [34], including 29 images of various objects and styles. While the original dataset provides prompts, these are generic image editing prompts. Since our work primarily focuses on re-rendering, we keep the editing prompts with style hints in the original dataset and added additional style prompts, resulting in a total of 8 prompts for each image. This benchmark assesses stylized re-rendering and structure preservation.
CARLA is an open-source driving simulator [8]. We collect 5.5 hours of driving videos from CARLA Town 4 using the simulator’s default autopilot. We then split the videos into 25 second clips and annotate a caption for each clip. For simulation enhancement, we use these captions combined with the style hint “A photorealistic video of driving”. We evaluate the effectiveness of sim-to-real transfer by testing the CARLA-trained planner on Waymo’s WOD-E2E [39] validation set.
@@ 4.2 Results
Figure 9: Video re-rendering results for simulation enhancement. From top to bottom are: input, Cosmos-Transfer2.5 [22], ours.
@@ Appendix A Additional Implementation Details
We implemented $\phi$-PD based on three models: SD 1.5 [29], FLUX-dev [18] and Wan2.2-14b [33]. We use the implementation of these models from DiffSynthStudio [^2] https://github.com/modelscope/DiffSynth-Studio This section describes the additional implementation details for the experiment with each model.
@@ A.1 Training and Inference Details
We start from the officially released checkpoints of SD 1.5 [29], FLUX-dev [18], and Wan2.2-14B [33], and finetune each model with phase-preserving noise. For SD 1.5, we experiment with both full finetuning and LoRA finetuning, while for FLUX-dev and Wan2.2-14B we use LoRA finetuning due to computational constraints. At inference time, for Wan2.2-14B we adopt the 4-step LoRA from LightX2V [^3] https://huggingface.co/lightx2v/Wan2.2-Lightning and apply it directly on top of our finetuned LoRA weights to accelerate sampling.